# A Simple and Effective Pruning Approach for Large Language Models

### v1

@@ Abstract
As their size increases, Large Languages Models (LLMs) are natural candidates for network pruning methods: approaches that drop a subset of network weights while striving to preserve performance. Existing methods, however, require either retraining, which is rarely affordable for billion-scale LLMs, or solving a weight reconstruction problem reliant on second-order information, which may also be computationally expensive. In this paper, we introduce a novel, straightforward yet effective pruning method, termed Wanda (Pruning by Weights and activations), designed to induce sparsity in pretrained LLMs. Motivated by the recent observation of emergent large magnitude features in LLMs, our approach prune weights with the smallest magnitudes multiplied by the corresponding input activations, on a per-output basis. Notably, Wanda requires no retraining or weight update, and the pruned LLM can be used as is. We conduct a thorough evaluation of our method on LLaMA across various language benchmarks. Wanda significantly outperforms the established baseline of magnitude pruning and competes favorably against recent methods involving intensive weight update. Code is available at https://github.com/locuslab/wanda.
@@ 1 Introduction
Large language models Touvron et al. (2023); Brown et al. (2020); OpenAI (2023); Hoffmann et al. (2022); Zhang et al. (2022); Scao et al. (2022) have recently reshaped the field of NLP with their remarkable performance across a range of complex language benchmarks Bommarito and Katz (2022); Wei et al. (2022b); Bubeck et al. (2023); Wei et al. (2022a). However, these models, with their billions of parameters, usually require significant computational resources. To democratize LLMs, considerable efforts have been taken to mitigate their high computational cost. Many of the notable advancements to date have centered on model quantization, a process where parameters are quantized into lower bit-level representations. The fast pace of LLM quantization research Dettmers et al. (2022); Xiao et al. (2023); Frantar et al. (2023); Yao et al. (2022); Dettmers and Zettlemoyer (2022); Xiuying et al. (2022); Lin et al. (2023); Ahmadian et al. (2023); Dettmers et al. (2023) has led to substantial resource savings for these models llama.cpp ; Sheng et al. (2023).
@@
Network pruning Hassibi et al. (1993); LeCun et al. (1989); Han et al. (2015); Liu et al. (2019); Jonathan and Michael (2019); Molchanov et al. (2017), on the other hand, shrinks network sizes by removing specific weights from the model – essentially setting them to zero. Along with quantization, it is often considered another popular approach for compressing neural networks. However, it has received relatively little focus in compressing LLMs. This seems to contradict the trend of model compression in the pre-LLM era, where both approaches have received large amounts of research effort. A quick review of existing pruning methods reveals a possible reason: they typically require retraining Han et al. (2015); Blalock et al. (2020), training from scratch Jonathan and Michael (2019); Renda et al. (2020); Louizos et al. (2018); Chen et al. (2020) or even an extensive iterative process Zhu and Gupta (2017); Gale et al. (2019); Li et al. (2020). The sheer computational resources required by LLMs limit these methods. A recent LLM pruning approach, SparseGPT Frantar and Alistarh (2023), does not require traditional retraining, but still demands a computationally intensive weight update process.
@@
The argument concerning the need for retraining and weight update does not fully capture the challenges of pruning LLMs. One might reasonably expect to obtain a fairly high-performing initialization point for retraining using existing popular pruning methods. However, a recent study Frantar and Alistarh (2023) finds that magnitude pruning Han et al. (2015); Zhu and Gupta (2017); Gale et al. (2019), a well-established pruning approach, fails dramatically on LLMs with relatively low levels of sparsity. Considering the past success of magnitude pruning on smaller networks, this result suggests that LLMs, despite having 100 to 1000 times more parameters, are substantially more difficult to prune directly.
@@
In this work, we address this challenge by introducing a straightforward and effective approach, termed Wanda (Pruning by Weights and activations). This technique successfully prunes LLMs to high degrees of sparsity without any need for modifying the remaining weights. This is inspired by insights from a recent study Dettmers et al. (2022) observing that a small subset of hidden state features are exceptionally large in magnitude, a property unique to LLMs. We find that augmenting the standard weight magnitude pruning metric with the input activations, is surprisingly effective as a measure for evaluating the weight importance. Specifically, we introduce a novel pruning metric, where each weight is evaluated by the product of its magnitude and the norm of the corresponding input activations, estimated using a small set of calibration data. Our method uses this metric to induce sparsity in pretrained LLMs by comparing weights locally within each output of linear layers and removing lower priority weights. Our approach is computationally efficient, able to be executed in a single forward pass, and requires minimal memory overhead.
@@
We empirically evaluate Wanda on LLaMA (Touvron et al., 2023), one of the most effective open-sourced LLM families. Our results demonstrate Wanda can find efficient sparse networks directly from pretrained LLMs, without any retraining or weight update. Wanda outperforms magnitude pruning by a large margin and also outperforms or matches the performance of a recently proposed method for pruning LLMs (Frantar and Alistarh, 2023), while requiring a lower computational cost. We hope our work serves as a baseline for future work in this area, and encourage further exploration in understanding sparsity in LLMs.
@@ 2 Preliminaries
Magnitude Pruning (Han et al., 2015, 2016) is a standard pruning technique to induce sparsity in neural networks. Different from structured pruning approaches Xia et al. (2022b); Liu et al. (2017), magnitude pruning removes individual weights based on their magnitudes, where weights with magnitudes below a certain threshold are removed. In practice, this threshold is typically determined by comparing weights globally Liu et al. (2019) or layer-wise Zhu and Gupta (2017). Despite its simplicity, magnitude pruning has been used to find extremely sparse networks Jonathan and Michael (2019); Zhu and Gupta (2017) and now stands out as a strong baseline approach Gale et al. (2019); Blalock et al. (2020); Hoefler et al. (2021) for neural network sparsification.
Emergent Large Magnitude Features is a property specific to Transformer Vaswani et al. (2017) based large language models. Once models reach a certain scale (in practice, around 6B parameters), a small set of hidden state features emerges with significantly larger magnitudes than the remaining ones. The existence of these outlier features in large language models is demonstrated in Dettmers et al. (2022). These outlier features exhibit several intriguing characteristics. First, they have very large magnitudes, about 100 times larger than typical hidden state values (Xiao et al., 2023). Second, they are usually sparse and exist in certain feature dimensions. Finally, these outlier features are essential for the predictive capability of LLMs: zeroing out these features at inference time results in significant degradation of language modeling performance (Dettmers et al., 2022).
3 Pruning by Weights and Activations
In this section, we motivate and describe our proposed pruning approach, Wanda (Pruning by Weights and activations), which consists of two simple but essential components. First, we propose a novel pruning metric that incorporates both weights and input activations into the computation of weight importance. Second, we compare weights on a per-output basis instead of across the whole layer, a factor that we find is crucial for pruning LLMs effectively. In Figure 1, we show an overview of our pruning approach.
We first start with a motivating example. Consider a neuron with two inputs and corresponding weights: , where . Now suppose the goal is to select one weight for removal while incurring less change on the output. The standard approach of magnitude pruning would always remove weight , which may be a good strategy if input features and have similar magnitudes. However, as recently observed in LLMs Dettmers et al. (2022), the two input features can differ significantly in scale. For instance, it is possible that , and as a result, . In this case, we should remove weight instead, because this clearly exerts a smaller influence on the neuron output than removing weight .
@@ 3 Pruning by Weights and Activations
Pruning Metric. Consider a fully connected layer with weight of shape . For language models, this linear layer takes in input activations with a shape of , where and are batch and sequence dimensions respectively. For each individual weight, we propose to evaluate its importance by the product of its magnitude and the corresponding input feature norm. Specifically, the score for the current weight is defined by:
@@
where represents the absolute value operator, evaluates the norm of th features aggregated across different tokens, and the final score is computed by the product of these two scalar values. We find that norm tends to work better than other norm functions in measuring activation magnitudes, for instance, and norm. We hypothesize this is likely because is generally a smoother metric Cortes et al. (2009); Lewkowycz and Gur-Ari (2020).
@@
This metric is interesting in several aspects. First, when the input channel of the considered weight has large magnitude features, the weight itself tends to be assigned a larger importance score despite potentially having a low magnitude. This tackles the problem we encounter in the motivating example. The effect can be seen in Figure 1, where weights corresponding to the large magnitude feature are more likely to be preserved with our approach Wanda. Second, its computation is straightforward. Once we obtain the norm vector of input feature activations, the weight importance can be calculated using an element-wise dot product. Last, empirical evidence from our experiments suggests that this metric is robust and can be easily estimated using a modest number of calibration samples (Section 4.5), without access to the original training data.
Pruning Granularity. To motivate the second aspect of our approach, we argue that in addition to the pruning metric, the pruning granularity, i.e. the set of weights among which to compare the importance, plays a significant role in pruning LLMs. Existing pruning methods typically employ a layer-wise Zhu and Gupta (2017) or global Liu et al. (2023); Sanh et al. (2020) pruning strategy, meaning that weights are compared either within each layer or across all layers in the network. This practice has been predominant in magnitude pruning approaches, often successfully leading to highly sparse networks Jonathan and Michael (2019); Gale et al. (2019).
@@
While the choice of pruning granularity has often been taken for granted in prior pruning approaches Jonathan and Michael (2019); Liu et al. (2023), we suggest that pruning LLMs could benefit from more local granularity levels. In our approach, we compare and remove weights on a per-output basis (per row in Figure 1), where weight importance scores are compared locally within each output neuron. Specifically, for a weight that connects input to output inside the linear layer, the group of comparison for this weight is defined as all weights connecting to output :
@@
Under this granularity, for a pre-defined sparsity ratio , we eliminate of the weights connected to each output. This practice may seem counter-intuitive, since we are basically pruning under a stricter sparsity pattern. However, we find that it is consistently better than layer-wise pruning for LLMs. A possible reason is that removing weights at a uniform ratio per output helps prevent imbalanced pruning across different output features.
@@
We find that for LLMs, removing weights at a per-output level not only enhances our proposed pruning metric but also the standard magnitude metric. To see if this holds true in general, we conduct additional experiments on pruning image classifiers. We observe no similar trend in image classification models, suggesting that observations regarding pruning per output on LLMs might be due to certain properties of the language modeling task. We hope this intriguing observation encourages practitioners to be more cautious in choosing the pruning granularity.
@@
Procedure. Wanda can be implemented and integrated seamlessly within a single forward pass of the LLM model, where feature norm statistics are estimated with a set of calibration data. We provide the PyTorch code of our approach in Algorithm 1. Given a pretrained LLM, we compute our pruning metric from the initial to the final layers of the network. After pruning a preceding layer, the subsequent layer receives updated input activations, based on which its pruning metric will be computed. A recent method for pruning LLMs, SparseGPT (Frantar and Alistarh, 2023), requires iterative weight update with a sophisticated procedure in the gradual pruning process, while Wanda requires no weight update. The sparse LLM after pruning is ready to use without further training or weight adjustment.
@@
Structured N:M Sparsity. While our method Wanda so far has been developed for unstructured sparsity, it can be easily extended to structured N:M sparsity (Zhou et al., 2021; Hubara et al., 2021; Pool and Yu, 2021). N:M sparsity requires that at most N out of every M contiguous weights to be non-zero. It can leverage NVIDIA’s sparse tensor cores (Mishra et al., 2021) to accelerate matrix multiplication in practice. Wanda can be naturally extended to structured N:M sparsity, where we only need to compare weights using the same metric among every M consecutive weights, for all weights connected to an output.
@@
Remark. We discuss Wanda’s connection with a few existing works. SparseGPT (Frantar and Alistarh, 2023) formalizes the problem of pruning LLMs by solving a local layer-wise reconstruction problem, where their pruning metric and weight update procedure is inspired from Optimal Brain Surgeon (OBS) (Hassibi et al., 1993). The pruning metric in SparseGPT is:
@@
Here in the denominator is the Hessian for the layer-wise reconstruction problem and is the Hessian dampening factor to avoid the collapse of inverse computation. With careful inspection, we observe that our metric in Equation 1 is similar to the above when is 0 and only the diagonal elements of the Hessian matrix are retained. This is because the diagonal of is , and the denominator in Equation 3 is now simplified to . The resulting metric is the square of our proposed metric. This simplification substantially reduces the required computation of weight importance, eliminating the need for computing any matrix inverses.
@@
In the 1980s, LeCun et al. (LeCun et al., 1989) have set up a pioneering framework for neural network pruning named Optimal Brain Damage (OBD). It uses second-order information but ignores off-diagonal elements in Hessians for faster approximation. Later, Optimal Brain Surgeon (OBS) (Hassibi et al., 1993) develops upon OBD partly by taking into account the off-diagonal elements. Wanda can be seen as a renaissance of the pioneering work of Optimal Brain Damage (OBD) (LeCun et al., 1989) – it may be viewed as applying a process similar to OBD to each neuron, with local output reconstruction as the objective function, whereas the original OBD uses the global training objective.
@@
| Method | Retraining | Weight Update | Calibration Data | Pruning Metric | Complexity |
| Magnitude | ✗ | ✗ | ✗ | ||
| SparseGPT | ✗ | ✓ | ✓ | ||
| Wanda (ours) | ✗ | ✗ | ✓ |
@@
A comparison of LLM pruning methods can be found in Table 1. Here the time complexity of computing the Wanda pruning metric is reduced compared with SparseGPT because it does not involve inverse computation. Overall, our method Wanda (Pruning by Weights and activations) has several attractive properties as an approach for pruning LLMs:
@@
Wanda can be done with a single forward pass of the LLM. At each layer, the pruned weights can be decided in one shot without an iterative procedure. In practice, Wanda can be 300 times faster in pruning LLMs compared with SparseGPT (Frantar and Alistarh, 2023).
@@
Unlike SparseGPT, our approach entails no weight updates on pruned networks, suggesting that LLMs have effective sparse sub-networks that are exact, instead of them merely existing in the neighborhood of the original weights.

### v3

@@ Abstract
As their size increases, Large Languages Models (LLMs) are natural candidates for network pruning methods: approaches that drop a subset of network weights while striving to preserve performance. Existing methods, however, require either retraining, which is rarely affordable for billion-scale LLMs, or solving a weight reconstruction problem reliant on second-order information, which may also be computationally expensive. In this paper, we introduce a novel, straightforward yet effective pruning method, termed Wanda (Pruning by Weights and activations), designed to induce sparsity in pretrained LLMs. Motivated by the recent observation of emergent large magnitude features in LLMs, our approach prunes weights with the smallest magnitudes multiplied by the corresponding input activations, on a per-output basis. Notably, Wanda requires no retraining or weight update, and the pruned LLM can be used as is. We conduct a thorough evaluation of our method Wanda on LLaMA and LLaMA-2 across various language benchmarks. Wanda significantly outperforms the established baseline of magnitude pruning and performs competitively against recent method involving intensive weight update. Code is available at https://github.com/locuslab/wanda.
@@ 1 Introduction
Large language models (Brown et al., 2020; OpenAI, 2023) have recently reshaped the field of NLP with their remarkable performance across a range of complex language benchmarks (Bommarito & Katz, 2022; Wei et al., 2022a; Bubeck et al., 2023). However, these models, with their billions of parameters, usually require significant computational resources. To democratize LLMs, considerable efforts have been taken to mitigate their high computational cost. Many of the notable advancements to date have centered on model quantization, a process where parameters are quantized into lower bit-level representations. The fast pace of LLM quantization research (Dettmers et al., 2022; Frantar et al., 2023a; Xiao et al., 2023; Ahmadian et al., 2023) has led to substantial resource savings for these models (Sheng et al., 2023; Lin et al., 2023).
@@
Network pruning (LeCun et al., 1989; Hassibi et al., 1993; Han et al., 2015), on the other hand, shrinks network sizes by removing specific weights from the model – essentially setting them to zero. Along with quantization, it is often considered another popular approach for compressing neural networks. However, it has received relatively little focus in compressing LLMs. This seems to contradict the trend of model compression in the pre-LLM era, where both approaches have received large amounts of research effort. A quick review of existing pruning methods reveals a possible reason: they typically require retraining (Liu et al., 2019; Blalock et al., 2020), training from random initializations (Zhu & Gupta, 2017; Louizos et al., 2018; Gale et al., 2019) or even an extensive iterative process (Frankle & Michael, 2019; Renda et al., 2020). The sheer amount of computational resources required by LLMs limits these methods. A recent LLM pruning approach, SparseGPT (Frantar & Alistarh, 2023), does not require traditional retraining, but still demands a computationally intensive weight update process.
@@
The argument concerning the need for retraining and weight update does not fully capture the challenges of pruning LLMs. One might reasonably expect to obtain a fairly high-performing initialization point for retraining using existing popular pruning methods. However, a recent study (Frantar & Alistarh, 2023) finds that magnitude pruning (Han et al., 2015), a well-established pruning approach, fails dramatically on LLMs even with relatively low levels of sparsity. Considering the past success of magnitude pruning on smaller networks, this result suggests that LLMs, despite having 100 to 1000 times more parameters, are substantially more difficult to prune directly.
@@
In this work, we address this challenge by introducing a straightforward and effective approach, termed Wanda (Pruning by Weights and activations). This technique successfully prunes LLMs to high degrees of sparsity without any need for modifying the remaining weights. We are motivated by an observation from a recent study (Dettmers et al., 2022), where a small subset of hidden state features are exceptionally large in magnitude, a property unique to LLMs. We find that augmenting the standard weight magnitude pruning metric with the input activations, is surprisingly effective as a measure for evaluating the weight importance. Specifically, we introduce a novel pruning metric, where each weight is evaluated by the product of its magnitude and the norm of the corresponding input activations, estimated using a small set of calibration data. Our method uses this metric to induce sparsity in pretrained LLMs by comparing weights locally within each output of linear layers and removing lower priority weights. Our approach is computationally efficient, able to be executed in a single forward pass, and requires minimal memory overhead.
@@
We empirically evaluate Wanda on the widely adopted LLaMA (Touvron et al., 2023a) and LLaMA-2 (Touvron et al., 2023b) model families. Our results demonstrate Wanda can find efficient sparse networks from pretrained LLMs, without any retraining or weight update. Our approach Wanda outperforms the standard magnitude pruning by a large margin and also competes favorably with the prior best LLM pruning method (Frantar & Alistarh, 2023), while requiring a lower computational cost. We hope our work serves as a baseline for future work in this area, and encourages further exploration in understanding sparsity in LLMs.
@@ 2 Preliminaries
Magnitude Pruning (Han et al., 2015) is a standard pruning technique to induce sparsity in neural networks. It removes individual weights based on their magnitudes, where weights with magnitudes below a certain threshold are removed. In practice, this threshold is typically determined by comparing weights locally within each layer or globally across the whole network. Despite its simplicity, magnitude pruning has been used to find extremely sparse networks (Frankle & Michael, 2019) and now stands out as a strong baseline approach (Blalock et al., 2020) for neural network sparsification.
Emergent Large Magnitude Features have been observed in Transformer-based large language models. Dettmers et al. (2022) discover that once LLMs reach a certain scale (in practice, around 6B parameters), a small set of hidden state features emerges with significantly larger magnitudes than the remaining ones. These outlier features exhibit several intriguing characteristics. First, they have very large magnitudes, about 100 times larger than typical hidden state values. Second, they are usually sparse and exist in certain feature dimensions. Finally, these outlier features are essential for the predictive capability of LLMs: zeroing out these features at inference time results in significant degradation of language modeling performance.
3 Wanda: Pruning by Weights and Activations
In this section, we motivate and describe our pruning method, Wanda (Pruning by Weights and activations), which consists of two simple but essential components. First, we propose a novel pruning metric that incorporates both weights and input activations into the computation of weight importance. Second, we compare weights on a per-output basis instead of across the whole layer, which we find is crucial for pruning LLMs effectively. An overview of Wanda is shown in Figure 1.
A Motivating Example. Consider a neuron with two inputs and corresponding weights: , where . Now suppose the goal is to select one weight for removal while incurring less change on the output. The standard approach of magnitude pruning would always remove weight , which may be a good strategy if input features and have similar magnitudes. However, as recently observed in LLMs (Dettmers et al., 2022), the two input features can differ significantly in scale. For instance, it is possible that , and as a result, . In this case, we should remove weight instead, because this removal clearly exerts a smaller influence on the neuron output than removing weight .
@@ 3 Wanda: Pruning by Weights and Activations
Pruning Metric. Consider a linear layer with weight of shape . For language models, this linear layer takes in input activations with a shape of , where and are batch and sequence dimensions respectively. For each individual weight, we propose to evaluate its importance by the product of its magnitude and the corresponding input feature norm. Specifically, the score for the current weight is defined by:
@@
where represents the absolute value operator, evaluates the norm of th features aggregated across different tokens, and the final score is computed by the product of these two scalar values. We find that norm tends to work better than other norm functions (e.g., and ) in measuring activation magnitudes. This is possibly because norm is generally a smoother metric.
@@
This metric is interesting in several aspects. First, when the input channel of the considered weight has large magnitude features, the weight itself tends to be assigned a larger importance score even if it has a low magnitude. This tackles the problem we encounter in the motivating example. The effect can be seen in Figure 1, where weights corresponding to the large magnitude feature are more likely to be preserved with Wanda. Second, its computation is straightforward. Once we obtain the norm vector of input feature activations, the weight importance can be calculated using an element-wise dot product. Last, we find empirically that this metric is robust and can be easily estimated using a modest number of calibration samples, without access to the original training data.
Comparison Group. Generally, in a pruning method, each weight is first assigned an importance score, such as the pruning metric we discussed above. These weights are then grouped into comparison groups where weights within each group are compared against one another. Within each comparison group, weights with lower importance scores are pruned. Most previous pruning methods default to comparing weights locally within each layer or globally across the whole network.
@@
While layer-wise and whole-network comparisons have been the popular options, we find that pruning LLMs could benefit from a more localized grouping. In our method, we compare and remove weights on a per-output basis (per row in Figure 1), where weight importance scores are compared locally within each output neuron. Specifically, for a weight that connects input to output inside the linear layer, we define the comparison group for this weight as all weights connecting to output :
@@
Under this comparison group, for a pre-defined sparsity ratio , we eliminate of the weights connected to each output. This practice may seem counter-intuitive, since we are basically pruning under a stricter sparsity pattern. However, we find that it is consistently better than layer-wise pruning for LLMs. Notably, this holds true not only for our proposed pruning metric (Equation 1) but also the standard magnitude metric. This shows that maintaining a balanced pruning ratio across output features is important for pruning LLMs effectively.
@@
To see if the superiority of pruning per output over per layer holds true in general, we conduct additional experiments on pruning image classifiers. However, we do not observe similar trend in image classification models, suggesting that our observations regarding pruning per output might be unique to LLMs. We hope this intriguing observation encourages practitioners to be more cautious in choosing the comparison group.
@@
Procedure. Wanda can be implemented and integrated seamlessly within a single forward pass of the LLM model, where feature norm statistics are estimated with a set of calibration data. We provide the PyTorch code of our approach in Algorithm 1. Given a pretrained LLM, we compute our pruning metric from the initial to the final layers of the network. After pruning a preceding layer, the subsequent layer receives updated input activations, obtained on the pruned weights of the previous layer. Then the pruning metrics are computed. A recent method for pruning LLMs, SparseGPT (Frantar & Alistarh, 2023), requires a sophisticated weight update procedure in an iterative pruning process, while Wanda does not induce any additional weight update.
@@
Structured N:M Sparsity. While Wanda so far has been developed for unstructured sparsity, it can be easily extended to structured N:M sparsity (Mishra et al., 2021). Structured N:M sparsity requires that at most N out of every M contiguous weights to be non-zero. It can leverage NVIDIA’s sparse tensor cores to accelerate matrix multiplication in practice. Wanda can be naturally extended to structured N:M sparsity, where we compare weights using the same metric among every M consecutive weights, for all weights connected to an output.
@@
Remark. We discuss the connection between Wanda and a few existing works. SparseGPT formalizes the problem of pruning LLMs by solving a local layer-wise reconstruction problem, where their pruning metric and weight update procedure is inspired from Optimal Brain Surgeon (OBS) (Hassibi et al., 1993). The pruning metric in SparseGPT is:
@@
Here in the denominator is the Hessian for the layer-wise reconstruction problem and is the Hessian dampening factor to avoid the collapse of inverse computation. With careful inspection, we observe that our metric in Equation 1 is similar to the above when is 0 and only the diagonal elements of the Hessian matrix are retained. Starting from the pruning metric in Equation 3, we show the exact reduction steps and corresponding reduction conditions as follows:
In the last reduction step, the diagonal of is , and thus the denominator can be simplified to . The resulting metric in Equation 4 is the square of our proposed metric. This simplification substantially reduces the required computation of weight importance, eliminating the need for computing any matrix inverses.
@@
In the 1980s, LeCun et al. (1989) have set up a pioneering framework for neural network pruning named Optimal Brain Damage (OBD). It uses second-order information without off-diagonal elements in Hessians for faster approximation. Later, Optimal Brain Surgeon (OBS) develops upon OBD partly by taking into account the off-diagonal elements. Wanda can be seen as a renaissance of OBD – it may be viewed as applying a process similar to OBD to each neuron, with local output reconstruction as the objective function, whereas the original OBD uses the global training objective. This is analogous to the relationship between SparseGPT and OBS.
@@
| Method | Weight Update | Calibration Data | Pruning Metric | Complexity |
| Magnitude | ✗ | ✗ | ||
| SparseGPT | ✓ | ✓ | ||
| Wanda | ✗ | ✓ |
@@
A comparison of LLM pruning methods can be found in Table 1. Computing the pruning metric of Wanda has a reduced time complexity compared to SparseGPT, because it does not involve inverse computation. Overall, our method Wanda (Pruning by Weights and activations) has several attractive properties as an approach for pruning LLMs:
@@
Wanda can be done with a single forward pass of the LLM. At each layer, the pruned weights can be decided in one shot without an iterative procedure. In practice, computing the pruning metric of Wanda can be 300 times faster in pruning LLMs compared with SparseGPT.
@@
Unlike SparseGPT, our approach entails no weight update on pruned networks, suggesting that LLMs have effective sparse sub-networks that are exact, instead of them merely existing in the neighborhood of the original weights.